OpenStudy (ashking1):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (ashking1):

answer choices 3.0° 6.0° -7.0° 16.0°

OpenStudy (perl):

Hint: $$u.v = |u| |v| \cos \theta$$

OpenStudy (ashking1):

so would i add the 2 numbers in u together and in v

OpenStudy (perl):

first evaluate the dot product we could solve for \( \theta \)

OpenStudy (ashking1):

im so lost in this question sorry. how would i begin by solving this

OpenStudy (perl):

@ashking we we can solve for theta \[ \Large{\sf u.v = |u| |v| \cos \theta \\ \quad \\ \frac{u.v}{|u| |v| } = \cos \theta \\ \quad \\ \theta = \cos \left( \frac{u.v}{|u| |v| } \right)^{-1} }\]

OpenStudy (ashking1):

alright so the equation is changed to solve the problem a theta is the degrees do i begin plugging in number

OpenStudy (perl):

Yes. I assume you know what ' u.v ' and |u| mean.

OpenStudy (ashking1):

alright so i am at 38/2sqrt365\[38/2 \sqrt365\]

OpenStudy (ashking1):

i got it its 6.0

OpenStudy (perl):

Nice work .